# Image Can Bring Your Memory Back: A Novel Multi-Modal Guided Attack against Image Generation Model Unlearning

### v1

@@ 5. Experiments
P4D-N 36.00 82.00 70.00 8.00 22.00 12.00 52.00 14.00 2.00 84.00 38.20
5.1. Experimental Setup
Datasets
We adopt the original text prompts provided by UnlearnDiffAtk (Zhang et al., 2024), which are derived from the I2P dataset (Schramowski et al., 2023) and ChatGPT (OpenAI, 2023). The reference images for both UnlearnDiffAtk and our proposed Recall are sourced from Flux-Uncensored-V2 (Flu, [n. d.]) (for Nudity, Church, and Parachute) and Stable Diffusion v2.1 (SDV, [n. d.]) (for Van Gogh). Details of these prompts and reference images are provided in Appendix B.1, Table 8.
IGMU Methods
We evaluate our approach across ten state-of-the-art IGMU techniques: ESD (Gandikota et al., 2023), FMN (Zhang et al., 2024b), SPM (Lyu et al., 2024), AdvUnlearn (Zhang et al., 2024), MACE (Lu et al., 2024), RECE (Gong et al., 2024), DoCo (Wu et al., 2025), Receler (Huang et al., 2024), ConceptPrune (Chavhan et al., 2024), and UCE (Gandikota et al., 2024). Details on model weights and training configurations are provided in Appendix B.2.
Baselines
We compare our proposed Recall against several representative attack baselines: Text-only, Image-only, Text & R_noise, Text & Image, P4D (Chin et al., 2024), and UnlearnDiffAtk (Zhang et al., 2024). Their detailed descriptions and implementation can be found in Appendix B.3.
Evaluation Metrics
To assess the effectiveness of our proposed attack, we employ task-specific deep learning-based detectors and classifiers to determine whether the target content has been successfully regenerated. These include the NudeNet detector (Praneeth, 2023), a ViT-based style classifier (Zhang et al., 2024), and an ImageNet-pretrained ResNet-50 (He et al., 2016). Detailed configurations of these models are provided in Appendix B.4. We report the Attack Success Rate (ASR, %) as the primary evaluation metric, defined as the percentage of generated images that contain the targeted concept. The average ASR across all ten unlearning techniques (denoted as Avg. ASR) is also computed to reflect overall robustness. To measure computational efficiency, we record the average attack time (in seconds) required to generate successful adversarial outputs. In addition, we compute the CLIP score (Hessel et al., 2021) between each generated image and its corresponding text prompt to evaluate semantic consistency, measuring how well the generated images align with the intended descriptions.
Implementation Details
Our Recall framework generates adversarial image prompts by leveraging reference-image guidance and performing perturbation optimization in the latent image space. The adversarial optimization is executed over 50 DDIM sampling steps, with 20 gradient ascent iterations per step using projected gradient descent (PGD (Madry et al., 2018)) with a step size of $\eta=1\mathrm{e}{-3}$ and a momentum coefficient of 0.9. An early stopping strategy is applied at timestep $t$ once the target content is successfully regenerated. The entire framework is implemented in PyTorch and evaluated on a high-performance server equipped with 8 NVIDIA H100 GPUs.
5.2. Attack Performance
@@
We comprehensively evaluate the effectiveness of our proposed Recall against several baseline attack methods across four representative unlearning tasks. The detailed experimental results, as summarized in Table 1, reveal several critical findings. ① Existing unlearning approaches do not really erase target concepts; notably, original textual or combined text-image prompts (reference image or randomly initialized) alone achieve substantial ASRs. For instance, combined text-image prompts yield an Avg. ASR exceeds 69.20% in the Parachute scenario. ② All baseline attack methods exhibit limited effectiveness when attacking adversarially enhanced unlearning strategies (e.g., AdvUnlearn and RECE), evidenced by their significantly lower ASRs. ③ In contrast, Recall consistently attains superior performance, achieving average ASRs ranging from 73.40% to 97.40% across diverse scenarios. Specifically, Recall outperforms UnlearnDiffAtk, a strong baseline, improving the average ASR by 16.90%, 0.20%, 11.00%, and 37.20% for Nudity, Van Gogh-style, Object-Church, and Object-Parachute, respectively. These results highlight the robustness and efficacy of Recall in regenerating targeted, presumably erased visual concepts.
Figure 3. Comparison of average attack time (in seconds) for different attack methods across three unlearning tasks: Nudity, Van Gogh-style, and Object-Parachute. The bar chart illustrates the attack efficiency of four attack approaches—P4D-K (blue), P4D-N (orange), UnlearnDiffAtk (red), and Recall (green)—against various unlearning techniques. A lower average attack time indicates more efficiency.
5.3. Attack Efficiency
@@
To assess the practical efficiency of Recall, we compare the average attack time with baseline methods, including P4D-K, P4D-N, and UnlearnDiffAtk. Figure 3 reports results across three representative tasks—Nudity, Van Gogh-style, and Object-Parachute—spanning multiple unlearning techniques. As shown, Recall achieves significantly lower attack time ($\sim$65s) compared to P4D-K ($\sim$380s), P4D-N ($\sim$340s), and UnlearnDiffAtk ($\sim$140s). This improvement stems from our efficient multi-modal optimization directly in the latent space, without relying on external classifiers or auxiliary diffusion models. Moreover, these efficiency gains align with our high attack success rates, highlighting that Recall is both effective and computationally lightweight. Notably, less robust unlearning methods (e.g., FMN, SPM) tend to require shorter attack durations, further illustrating their susceptibility  [^1] We exclude cases where initial prompts alone succeed, focusing on instances requiring iterative optimization..
5.4. Semantic Alignment Analysis
We assess the semantic consistency between regenerated images and their corresponding text prompts using the CLIP score. Table 2 presents the average CLIP scores for three attack methods—P4D, UnlearnDiffAtk, and our proposed Recall —evaluated across six unlearning techniques (ESD, MACE, RECE, UCE, Receler, and DoCo) and four aforementioned representative unlearning tasks.
As shown in Table 2, Recall consistently outperforms baseline methods, achieving the highest CLIP scores across all tasks and unlearning settings. Notably, Recall attains an average CLIP score of 30.28, surpassing UnlearnDiffAtk (28.00) and P4D (25.00). These results indicate that text-based methods, which perturb original prompts, often degrade semantic coherence. In contrast, our multi-modal adversarial framework preserves the textual intent and introduces perturbations solely through the image modality, yielding superior semantic alignment.
Table 2. Comparison of CLIP scores (higher is better) for images regenerated by existing attacks (P4D, UnlearnDiffAtk, and our proposed Recall) against various unlearning methods across four unlearning tasks. The best-performing attack method for each scenario is highlighted in bold.
@@ 5.4. Semantic Alignment Analysis
Table 3. Generated images under different attacks for MACE and RECE across different unlearning tasks.
@@
5.5. Visualization
Table 3 presents a qualitative comparison of regenerated images under four representative unlearning scenarios—Nudity, Van Gogh-style, Object-Church, and Object-Parachute—against two prominent unlearning techniques, MACE and RECE. Additional details, including text prompts, adversarial inputs, guidance scales, and seeds, are provided in Appendix D, Table 10.
Rows 3–6 show that neither original prompts nor their combination with random or reference images effectively bypass the safety filters of MACE and RECE. Although image-only settings perform better on object-centric tasks, they often lack semantic alignment and diversity. Incorporating text and reference images yields limited improvements and frequently fails to recover the erased concepts. The subsequent rows (7-9) show the generated images by P4D, UnlearnDiffAtk, and our proposed Recall. Recall consistently induces unlearned models to regenerate forgotten content with high semantic fidelity, whereas baseline methods frequently produce incomplete or inconsistent results. For Nudity, Recall reliably reconstructs sensitive visual features, outperforming P4D’s vague outputs. In Van Gogh-style, our method captures distinct artistic traits, unlike P4D’s distorted textures and UnlearnDiffAtk’s partial recovery. In Object-Parachute, Recall robustly restores the intended object, while others fail to ensure visual or semantic integrity. These findings underscore the limitations of existing text-based attacks and expose critical vulnerabilities in current unlearning strategies.
5.6. Ablation Study
In this section, we conduct ablation studies to systematically examine the generalizability of our method as well as the impact of important design choices and key hyperparameters. We first analyze the generalizability of our attack to variations in reference image selection, generation diversity, and model versions. We then investigate how different guidance strategies and optimization parameters affect the overall attack performance and semantic consistency. The following subsections provide detailed analyses and empirical results for each aspect.
5.6.1. Generalizability
Reference Independence.
To evaluate the robustness of our attack to the selection of reference images, we conducted experiments using three additional three different references ( $R_{1}-R_{3}$ shown in Appendix E.1 Figure 4) random download from the Internet, i.e., Nudity and Object-Church, respectively. As shown in Table 4, both the attack success rate (ASR,%) and diversity metrics (LPIPS (Zhang et al., 2018), Inception Score (IS) (Salimans et al., 2016), higher values indicate better diversity) remain consistently high across different choices of references, provided that the reference contains representative information for the target concept. This demonstrates that our method is robust to the reference source and does not depend on a specific image for effectiveness.

### v2

@@ 5.1 Experimental Setup
5.2 Attack Performance
Figure 3: Generated images under different attacks. Rows (top to bottom): Nudity, Van Gogh, Church, and Parachute.
@@ 5.2 Attack Performance
We comprehensively evaluate the effectiveness of Recall against several baseline attack methods across four representative unlearning tasks. The detailed experimental results, as summarized in Table 1, reveal several critical findings. ① Existing unlearning approaches fail to fully erase target concepts; notably, original textual or combined text-image prompts (reference image or randomly initialized) alone achieve substantial ASRs. For instance, combined text-image prompts yield an Avg. ASR exceeds 69.20% in the Parachute task. ② All baseline attack methods exhibit limited effectiveness when attacking adversarially enhanced unlearning strategies (e.g., AdvUnlearn and RECE), evidenced by their significantly lower ASRs. ③ In contrast, Recall consistently attains superior performance, achieving average ASRs ranging from 73.40% to 97.40% across diverse scenarios. Specifically, Recall outperforms UnlearnDiffAtk, a strong baseline, improving the average ASR by 16.90%, 0.20%, 11.00%, and 37.20% for four tasks. These results highlight the robustness and efficacy of Recall in regenerating targeted, presumably erased visual concepts.
In addition, qualitative generation results on MACE in Figure 3 (Complete results in Appendix D Table 4) and visual cases (in Appendix F.2 Figure 7) show that Recall consistently surpasses existing baselines in recovering erased concepts across a variety of unlearning scenarios, yielding highly diverse outputs.
5.3 Attack Efficiency
Figure 4: Comparison of average attack time for different attack methods for Nudity task.
@@ 5.3 Attack Efficiency
To assess the practical efficiency of Recall, we compare the average attack time (in seconds) needed by Recall with various baselines, a lower average attack time indicates higher efficiency. Figure 4 reports results for Nudity task (attack time for more tasks can be found in Appendix E). As shown, Recall achieves significantly lower attack time ($\sim$64s) compared to P4D-N ($\sim$238s), UnlearnDiffAtk ($\sim$232s) and WACE-C ($\sim$243.15s). This improvement stems from our efficient multi-modal optimization directly in the latent space and these efficiency gains align with our high attack success rates, highlighting that Recall is both effective and computationally lightweight. Notably, less robust unlearning methods (e.g., FMN, SPM) tend to require shorter attack durations, further illustrating their susceptibility.
5.4 Semantic Alignment
Table 2: CLIP Scores ($\uparrow$) among baselines and Recall.
@@ 5.4 Semantic Alignment
The original SD model achieves CLIP Scores of 31.05, 33.98, 30.75, and 31.56 on the four tasks, respectively.
We assess the semantic consistency between regenerated images and their corresponding text prompts using the CLIP Score. Table 2 presents the average CLIP Scores for four attack methods, P4D, CCE, UnlearnDiffAtk, WACE, and our proposed Recall, evaluated across six unlearning techniques and four aforementioned representative unlearning tasks.
As shown in Table 2, Recall consistently outperforms baseline methods, achieving the highest CLIP Scores across all tasks and unlearning settings. Notably, Recall attains an average CLIP Score of 30.28, surpassing UnlearnDiffAtk (28.00), P4D (25.00), CCE (20.01) and WACE (19.89). These results indicate that text-based methods, which perturb original prompts, often degrade semantic coherence. In contrast, our multi-modal adversarial framework preserves the textual intent and introduces perturbations solely through the image modality, yielding superior semantic alignment.
5.5 Generalizability
Reference Independence. We assess the robustness of Recall to reference image selection using three additional references ($R_{1}$$R_{3}$, see Appendix F.1, Figure 6). These additional references are randomly downloaded from the Internet for the Nudity and Object-Church tasks; $R_{\text{org}}$ is the default reference in our core experiments, while $R_{1}$$R_{3}$ are used only for this ablation. As reported in Appendix F.1, Table 5, both attack and diversity metrics remain consistently high provided the reference is representative, demonstrating that Recall does not depend on any specific image.
Generation Diversity. We quantitatively compare image-only, text-only, and Recall. As detailed in Appendix F.2, Recall achieves substantially greater diversity than image-only baselines and matches the performance of text-only approaches, indicating that it recovers the original concept distribution rather than simply transforming reference images.
Model Version Independence. We further evaluate Recall on unlearned models based on SD 2.0 and SD 2.1, in addition to SD 1.4. As shown in Appendix F.3, Recall consistently maintains high effectiveness across all versions, confirming its robustness and generalizability to more advanced diffusion architectures.
5.6 Ablation Study
We conduct ablation studies to systematically evaluate the impact of key strategies and hyperparameters on the performance of the Recall framework.
Strategies. We analyze three core strategies: multi-modal guidance, noise initialization, and periodic integration.
  • Multi-modal Guidance. We compare Text-only, Image-only, Text & R_noise, Text & Image, and our Text & Adversarial Image approaches. Results in Sections 5.2 and Appendix D show that combining textual prompts with adversarial image optimization substantially improves both attack performance and semantic consistency.
  • Noise Initialization. Noise initialization significantly enhances both diversity and semantic alignment of generated images, as demonstrated by consistently higher LPIPS, IS, and CLIP Scores across tasks (Appendix G.1, Figure 9).
  • Periodic Integration. Periodically integrating $z_{ref}$ into $z_{adv}$ further improves attack performance, efficiency, and the diversity of generated images when the attack succeeds (Appendix G.4, Figure 12). We therefore adopt this strategy with $epoch_{\text{interval}}=5$ and $\gamma=0.05$.
Parameters. We investigate the sensitivity to two critical optimization parameters:
  • Step Size ($\eta$). Reducing $\eta$ from 0.1 to 0.001 steadily increases ASR, with $\eta=0.001$ yielding optimal performance. Further reduction impairs effectiveness due to insufficient updates (Appendix G.1, Figure 9).
  • Initial Balancing Factor ($\lambda$). Increasing $\lambda$ improves ASR until saturation. Semantic alignment (CLIP Score) peaks at $\lambda=0.25$ and then declines, at the same time, reaching a good tradeoff between the ASR and attack time, indicating a trade-off between attack strength and semantic consistency. We set $\lambda=0.25$ as the default (Appendix G.2, Figure 10).
6 Conclusion
We present Recall, a multi-modal adversarial framework for auditing concept unlearning in multi-modal conditioning IGMs. Distinct from previous text-based approaches, Recall leverages adversarially optimized image prompts together with the original textual inputs to induce unlearned IGMs to recover previously erased visual concepts. Extensive experiments across ten SOTA unlearning techniques and diverse tasks show that current pipelines remain vulnerable to multi-modal guided adversarial inputs. Beyond functioning as an attack, Recall provides an efficient auditing mechanism for model owners with full access to verify the robustness of their unlearning procedures prior to deployment, thereby informing the design of stronger, verifiable unlearning defenses.
Future Work. Future work will (i) extend Recall to black-box and transfer-based settings to enable third-party robustness auditing without parameter access; (ii) evaluate generalizability across broader generative architectures and training regimes; and (iii) investigate defense-aware and certifiable unlearning strategies that are resilient to multi-modal adversarial threats, including extensions to video and large multi-modal models.
Acknowledgments
This research is supported by A*STAR, CISCO Systems (USA) Pte.Ltd and National University of Singapore under its Cisco-NUS Accelerated Digital Economy Corporate Laboratory (Award I21001E0002).
Ethics Statement
This work evaluates vulnerabilities in concept-unlearned diffusion models using synthetic or publicly available data. We do not release explicit imagery; figures are masked where necessary. All experiments are conducted strictly for safety auditing and research purposes. We did not conduct any user studies or collect personal data.
Reproducibility Statement
We provide text prompts, reference images, random seeds, and hyperparameters in our publicly available codebase [^2] https://github.com/ryliu68/RECALL, and we summarize them in Appendix C (Table 3 and Table 7). We fix a global random seed to 2025. For each task, we use the task-specific image-generation seeds specified in our configuration (e.g., those used for the Nudity task on the I2P dataset), ensuring that all reported results are reproducible.
Appendix A Appendix Overview
This appendix provides supplementary material omitted from the main paper due to space constraints. Specifically, it includes:
  • Section B: Complete algorithmic procedure for the proposed Recall framework.
  • Section C: Detailed experimental setup, including datasets, unlearned IGMs, baseline methods, and evaluation metrics.
  • Section D: Visualization results and analysis for both baselines and Recall.
  • Section E: Comprehensive results on attack efficiency.
  • Section F: Detailed results on the generalizability of Recall, including reference image independence, generation diversity, and model version independence.
  • Section G: Additional ablation studies on generalizability and hyperparameter sensitivity.
Appendix B Algorithm
We list the Recall pipeline in Algorithm 1, which allows readers to re-implement our method step-by-step.
Algorithm 1 Recall
1:Input: Reference image $P_{\text{ref}}$, initial image $P_{\text{image}}^{\text{init}}$, text prompt $P_{\text{text}}$, diffusion model $\mathcal{G}_{u}$ (with U-Net $\mathcal{F}_{\theta}$, text encoder $\mathcal{E}_{t}$, image encoder $\mathcal{E}_{i}$, image decoder $\mathcal{D}_{i}$), hyperparameters $\lambda$, $\gamma$, $\eta$, $\beta$, number of DDIM steps $T$, PGD iterations NN
2:Output: Image $I^{*}$ with target content $t$ // Adversarial image optimization; Stage I, Stage II
[…]
8:for $t=T,T{-}1,\dots,1$ do
9:  $z_{\text{ref},t}\leftarrow\sqrt{\bar{\alpha}_{t}}z_{\text{ref}}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$, $\epsilon_{t}\sim\mathcal{N}(0,I)$
10:  $z_{\text{adv},t}\leftarrow\sqrt{\bar{\alpha}_{t}}z_{\text{adv}}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$
11:  $\hat{\epsilon}_{\text{ref}}\leftarrow\mathcal{F}_{\theta}(z_{\text{ref},t},t,h_{t})$
12:  $\hat{\epsilon}_{\text{adv}}\leftarrow\mathcal{F}_{\theta}(z_{\text{adv},t},t,h_{t})$
13:  $\mathcal{L}_{\text{adv}}\leftarrow\|\hat{\epsilon}_{\text{ref}}-\hat{\epsilon}_{\text{adv}}\|_{2}^{2}$
14:  Compute $\nabla_{z_{\text{adv}}}\mathcal{L}_{\text{adv}}$
15:  for $i=1$ to NN do
16:   $v_{i}\leftarrow\beta\cdot v_{i-1}+\nabla_{z_{\text{adv}}}\mathcal{L}_{\text{adv}}/(\|\nabla_{z_{\text{adv}}}\mathcal{L}_{\text{adv}}\|_{1}+\omega)$
17:   $z_{\text{adv}}\leftarrow z_{\text{adv}}+\eta\cdot\text{sign}(v_{i})$
18:  end for
19:  if $t\bmod\text{epoch}_{\text{interval}}=0$ then
20:   $z_{\text{adv}}\leftarrow z_{\text{adv}}+\gamma\cdot z_{\text{ref}}$
21:  end if
22:end for
[…]
25:Sample noise $\delta^{\prime}\sim\mathcal{N}(0,I)$
[…]
27:for $t=T,T-1,\dots,1$ do
28:  $\hat{\varepsilon}_{t}\leftarrow\mathcal{F}_{\theta}\bigl(z_{t},t,h_{t}\bigr)$
29:  $z_{t-1}\leftarrow\textit{SchedulerStep}\bigl(\hat{\varepsilon}_{t},t,z_{t}\bigr)$
30:end for
[…]
Appendix C Experimental Setup
C.1 Datasets
We evaluate our method on four unlearning tasks: 1) Nudity, 2) Van Gogh-style, 3) Object-Church, and 4) Object-Parachute to ensure a thorough examination of unlearned models’ vulnerabilities. Since multi-modal image generation consumes both text and image, we first collect a reference image with the sensitive content, note that these reference image can be soured from any place as it contain the target content, such as generated by generative models, internet, or make it with other tools. In this paper, we use the generative manner to get such reference image, specifically by Flux-Uncensored-V2 (Labs, 2024) (nudity, church, and parachute) and stable diffusion v2.1 (AI, 2024) (van Gogh) with a given text prompt for each task (as shown in Table 3), where the used text prompt are to be used for attacking. We then adopted the text prompts used in UnlearnDiffAtk (Zhang et al., 2024d) as the text prompts for each task, the details of these prompts are as follows:
  • Nudity: The dataset for this task are I2P, MMA, and ART. Inappropriate Image Prompts (I2P) dataset (Schramowski et al., 2023) is involved, which contains a diverse set of prompts leading to unsafe or harmful content generation, including nudity, we use the 142 nudity related prompts filtered by UnlearnDiffAtk (Zhang et al., 2024d). MMA (Yang et al., 2024) is 1000 adversarial optimized text prompts used to attack safe-checker of stable diffusion models. We also adopt the benign red-teaming dataset ART (Li et al., 2024a), which is automatically collected by ART framework from the Lexica gallery and focuses on benign prompts that still trigger harmful generations in text-to-image models. In addition to the above datasets, we also adopt the benign red-teaming benchmark introduced by ART. (Li et al., 2024a), which are collected (safe prompt, unsafe image) pairs from the Lexica gallery and focuses on benign prompts that still trigger harmful generations in text-to-image models. We use nudity-related prompts in our experiments.
  • Van Gogh-style: The prompts are artistic-painting prompts introduced in ESD (Gandikota et al., 2023), the number of prompts is 50.
  • Object-Church and Object-Parachute: The prompts are generated by GPT-4 (OpenAI, 2023b), and the number of prompts is 50 related to church and parachute, respectively.
Table 3: The reference image aligns with the text prompt and seed for the image-dependent attack method.
Task Nudity Van Gogh-style Object-Church Object-Parachute
Seed 6053 10710 33731 55254
C.2 Baselines
  • Text & Image: Both the original text prompt and a semantically relevant reference image containing the erased concept are provided as multi-modal inputs to the unlearned image generation models. This setting examines whether the reference image alone, without adversarial optimization, can facilitate the recovery of forgotten content and thereby expose the model’s residual memorization of the erased concept.
  • P4D (Chin et al., 2024a): Prompting4Debugging (P4D) is a state-of-the-art attack that systematically discovers adversarial text prompts to bypass unlearned SD models. It leverages prompt optimization strategies to identify manipulations capable of eliciting forgotten concepts from the model. We report the results of P4D-K and P4D-N in this part simultaneously. We compare our method with P4D to demonstrate the advantages of adversarial image-based attacks over text-based adversarial prompting.
  • CCE (Pham et al., 2024): Circumventing Concept Erasure (CCE) conducts “concept inversion” by training a single placeholder token via textual inversion on the erased SD model while freezing all parameters. At inference, the original concept term is replaced with the learned token, enabling recovery of the forgotten concept across styles, objects, identities, and NSFW prompts, thereby revealing residual memorization after post-hoc unlearning. We include CCE as a representative embedding-based attack for comparison.
  • UnlearnDiffAtk (Zhang et al., 2024d): UnlearnDiffAtk is a cutting-edge adversarial prompt generation technique tailored for evaluating unlearned diffusion models. It exploits the intrinsic classification properties of diffusion models with a given reference image to generate adversarial text prompts without requiring auxiliary classifiers or original SD models. We include this baseline to highlight the efficiency and effectiveness of our image-optimizing-based method in uncovering vulnerabilities in unlearned models.
  • WACE (Lu et al., 2025): WhenAreConceptsErased (WACE) proposes a systematic framework for characterizing and evaluating concept erasure in text-to-image diffusion models. It distinguishes between guidance-based avoidance and destruction-based removal, and introduces a multi-perspective probing suite that includes NoiseBasedProbe (WACE-N) and classifier-guided NoiseBasedProbe (WACE-C) to make the unlearned model regenerate the content assumed to be removed.
P4D (with its variants P4D-K and P4D-N) (Chin et al., 2024a) and UnlearnDiffAtk (Zhang et al., 2024d) are text-optimization-based attack methods. CCE injects a learned placeholder token into the unlearned model’s text embedding. WACE (with its varints WACE-N and WACE-C) is a noise-based attack method. And we use their officially released code (P4D [^3] https://github.com/joycenerd/P4D, CCE [^4] https://github.com/NYU-DICE-Lab/circumventing-concept-erasure, UnlearnDiffAtk [^5] https://github.com/OPTML-Group/Diffusion-MU-Attack, WACE [^6] https://github.com/kevinlu4588/WhenAreConceptsErased) with default configurations to extend the attack.
C.3 Evaluation Metrics
Attack performance.
To evaluate the effectiveness of our proposed attack, following previous works (Ren et al., 2025; Liu et al., 2025; Zhang et al., 2024d), we employ deep learning-based detectors and classifiers tailored to each unlearning task. For the Nudity task, we adopt the NudeNet detector (Praneeth, 2023) with a detection threshold $\tau=0.45$ to identify explicit anatomical features. Following standard protocol (Ren et al., 2025; Liu et al., 2025), an attack is considered successful if any of the following sensitive labels is detected from the generate image: MALE_BREAST_EXPOSED, MALE_GENITALIA_EXPOSED, FEMALE_BREAST_EXPOSED, FEMALE_GENITALIA_EXPOSED, BUTTOCKS_EXPOSED, or ANUS_EXPOSED. For the Van Gogh-style task, we use a ViT-based (Dosovitskiy et al., 2021) style classifier pretrained on ImageNet and fine-tuned on the WikiArt dataset (Saleh and Elgammal, 2015). Following (Zhang et al., 2024d), we use the top-10 predictions to determine whether the generated images exhibit Van Gogh’s characteristic artistic features. For the object-centric tasks, Object-Church and Object-Parachute, we use a ResNet-50 classifier pretrained on ImageNet to determine whether the corresponding object is present in the generated image using the top-1 prediction.
Semantic alignment.
For the alignment between the given text prompts and the images generated by successful attacks, we report the CLIP Score computed between each prompt–image pair and then averaged across prompts; higher values indicate stronger text–image consistency.
Image diversity.
To assess the diversity of images produced by successful attacks, we adopt three complementary metrics: LPIPS (Zhang et al., 2018), Inception Score (IS) (Salimans et al., 2016), and a DINO-based feature distance (Oquab et al., 2024).
Appendix D Visualization
@@
Table 4: Generated images under different attacks for MACE and RECE across different unlearning tasks.
@@
CCE [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
WACE [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 4 presents a qualitative comparison of regenerated images under four representative unlearning scenarios, i.e., Nudity, Van Gogh-style, Object-Church, and Object-Parachute, for the unlearning techniques MACE and RECE. Rows 3–6 illustrate that neither original prompts nor their combination with random or reference images effectively bypass the safety filters. While image-only settings perform somewhat better on object-centric tasks, they often lack semantic alignment and diversity; combining text and reference images yields only limited improvements.
The subsequent rows show results from P4D (Chin et al., 2024a), CCE (Pham et al., 2024), UnlearnDiffAtk (Zhang et al., 2024d), WACE (Lu et al., 2025), and our proposed Recall. Notably, baselines such as P4D and UnlearnDiffAtk typically require heavily modifying the input text to bypass unlearning, which can restore content but often at the cost of semantic fidelity—especially evident in the Nudity and Van Gogh-style scenarios. In contrast, Recall maintains the original prompt unchanged, leveraging adversarial image guidance to bypass unlearning while preserving strong semantic alignment.
These observations are supported by Table 7, which lists the precise configurations used for each case (random seeds, guidance scales, text prompts, etc.). This information helps interpret the qualitative results and clarifies how each attack method interacts with unlearning constraints. Overall, Recall consistently induces unlearned models to regenerate forgotten content with high semantic fidelity, outperforming existing baselines in both visual quality and semantic coherence.
Appendix E Attack Efficiency
Figure 5: Comparison of average attack time (in seconds) for different attack methods across three unlearning tasks. The bar chart illustrates the attack efficiency of four attack approaches—P4D-N (blue), UnlearnDiffAtk (orange), WACE-C (red), and Recall (green)—against various unlearning techniques. A lower average attack time indicates higher efficiency.
To quantitatively assess the practical advantage of the proposed Recall w.r.t computational efficiency, we evaluate and compare the average attack durations [^7] It is worth noting that we exclude cases where the initial prompts alone suffice to trigger successful attacks and consider only those instances where optimization is necessary for success. across various adversarial methods [^8] We omit CCE because it injects a learned placeholder token directly into the model’s text embedding by finetuning the model via textual inversion.. Figure 5 illustrates the average attack time (in seconds) [^9] Here, an attack time of 0 indicates that the attack succeeds on fewer than five prompts (out of all evaluated prompts) within the budget; thus, 0 denotes failure to meet our minimum-success threshold rather than negligible runtime. required by Recall and baseline methods, including P4D-N, UnlearnDiffAtk, and WACE-C against multiple unlearning techniques across three representative unlearning scenarios: Nudity, Van Gogh-style, and Object-Parachute.
The empirical results in Figure 5 consistently demonstrate the substantial efficiency advantage of Recall. Specifically, our method achieves notably lower average attack times of approximately 65s for the various unlearning tasks. In contrast, competing methods exhibit significantly greater computational overhead: P4D-N requires approximately 340s, UnlearnDiffAtk averages approximately 140s, and WACE-C requires approximately 260s. This considerable efficiency improvement can be attributed primarily to our multi-modal guided optimization approach conducted entirely in image latent space and eliminates reliance on external classifiers or auxiliary diffusion models.
Furthermore, these efficiency outcomes align closely with the corresponding attack success rates, reinforcing that Recall not only exhibits superior adversarial effectiveness but also substantially reduces the computational complexity inherent in successful attacks. Additionally, we observe that unlearning techniques with comparatively lower robustness, such as FMN and SPM, inherently require shorter attack durations, underscoring their heightened vulnerability in realistic adversarial scenarios.
Appendix F Generalizability
F.1 Reference Independence
Figure 6: Reference images used in our experiments. The top row corresponds to Nudity task, and the bottom shows Church task.
We put the additional reference images in Figure 6, which were randomly downloaded from the Internet for the Nudity and Object-Church tasks. $R_{\text{org}}$ is the main reference image used in the core experiments, while $R_{1}$, $R_{2}$, and $R_{3}$ are additional references introduced in the ablation study to assess the robustness and generalizability of our attack. The experimental results (Table 5) demonstrate that our Recall does not rely on any specific reference image. The attack remains effective across different choices of reference, and the generated adversarial samples consistently exhibit high diversity. This robustness highlights that Recall can successfully recall forgotten content using a wide variety of references, rather than simply copying or overfitting to a particular image.
Table 5 shows the attack performance and the generated images’ diversity with attack success rate (ASR,%) and diversity metrics (LPIPS, IS, and DINO). These results confirm that Recall does not rely on any particular reference image; across diverse reference sources, it achieves comparable attack performance while maintaining high diversity in the successfully generated images.